# On the universality of star formation efficiency in galaxies

### v1

@@ 1 Introduction
On the kiloparsec and larger scales galaxies also form stars inefficiently, which is partly related to the local efficiency per free-fall time (e.g., Krumholz et al., 2012; Semenov et al., 2018). The connection between efficiency in star-forming regions and star formation rate measured on large scales arises because the local inefficiency of star-forming regions and their short lifetimes result in only a small fraction of gas being converted into stars in each star-forming region. Therefore, a given gas parcel has to go through a large number of star formation-dispersal cycles, each taking $\approx 50-100$ Myr as ISM atoms make their way into a star-forming region (Semenov et al., 2017).
@@ 2 Simulations
We use a suite of simulations of an isolated disk galaxy adopting the initial conditions similar to the properties of the NGC 300 galaxy (Semenov et al., 2021). These simulations are run using the ART N-body$+$hydrodynamics code (Kravtsov, 1999; Kravtsov et al., 2002; Rudd et al., 2008; Gnedin & Kravtsov, 2011) and include self-consistent modeling of radiative transfer (RT; Gnedin 2014), non-equilibrium chemistry of hydrogen, helium, and molecular hydrogen, as well as realistic prescriptions for star formation and feedback. The latter is sufficiently realistic to replicate detailed properties of star forming regions in NGC 300 (Semenov et al., 2021).
@@
where the total sub-grid velocity dispersion, $\sigma_{\mathrm{tot}}=\sqrt{\sigma_{\mathrm{t}}^{2}+c_{\mathrm{s}}^{2}}$, includes contributions of both turbulence and the thermal energy of the gas. Thus, the local value of $\epsilon_{\rm ff}$ in simulation cells evolves with the evolution of the interstellar medium, based on its density and the level of turbulence on small scales. The latter is modeled using an explicit subgrid turbulence model (Semenov et al., 2016).
@@ 3.1 Star formation efficiency as a function of metallicity on $\sim 10$ pc scale
We note these characteristic efficiency values are not a result of tuning but arise naturally in simulations (see Section 4 for discussion). Furthermore, the top panels in Figures 3 and Figure 4 show that the characteristic $\epsilon_{\mathrm{ff}}$ range does not depend on gas metallicity and depends only very weakly on free-space [^1] Free-space UV flux is the flux at a given location not attenuated by local extinction. Krumholz et al. (2009), for example, use free-space flux to mean the flux incident on molecular clouds. In our simulations, the radiative transfer calculations do not include absorption by H2 lines and thus do not model radiation field self-consistently inside molecular-rich regions and have to rely on the subgrid model. In this case, the free-space flux is the flux returned by the radiative transfer solver and has the physical meaning of the incident field on the molecular gas. UV flux, $U_{\mathrm{MW}}$, normalized by the Milky Way value at 1000 Å (Draine, 1978; Mathis et al., 1983).
Remarkably, while the amount of star-forming gas decreases with decreasing metallicity due to increasingly inefficient cooling (see Polzin et al., 2024, for a detailed discussion), the range of $\epsilon_{\rm ff}$ remains the same for all simulated metallicities. In what follows we show that $\epsilon_{\rm ff}$ also has no or very weak dependence on the averaging scale and other properties of the local galactic environment.
@@ 3.2 Star formation efficiency as a function of local environment properties on 500 pc scale
We examine the star formation efficiency per free-fall time ($\epsilon_{\mathrm{ff}}$), the free-fall time ($t_{\mathrm{ff}}$), and the depletion time ($t_{\mathrm{ff}}/\epsilon_{\mathrm{ff}}$), averaged in patches of 500 pc size, as a function of metallicity, UV flux, and baryon (stars and gas) surface density. Namely, we compute the weighted medians [^2] We define weighted quantiles by the value corresponding to the percentile of the cumulative distribution of weights ordered by the product of the values and the weights. of these quantities using the normalized local star formation rate in each cell as a weight. We also only use the cells with $\alpha_{\mathrm{vir}}\leq 15$, which contribute the bulk of the total star formation in each simulated galaxy. Note that we actually compute the SFR-weighted averages of $1/t_{\mathrm{ff}}$ and $\epsilon_{\mathrm{ff}}/t_{\mathrm{ff}}$ because the star formation rate is proportional to these quantities. However, we plot their inverses $t_{\mathrm{ff}}$ and $t_{\mathrm{ff}}/\epsilon_{\mathrm{ff}}$ the free-fall time and the depletion time values are more easily interpretable.
@@ 4 Discussion
As noted above, the characteristic values of $\epsilon_{\mathrm{ff}}$in our simulations arise naturally from the specific star-formation prescription based on the simulations of star formation in turbulent molecular clouds (see eq. 1 above; Padoan et al., 2012). The efficiency in this simulation-calibrated prescription is an exponential function of $\sqrt{\alpha_{\rm vir}}$, where $\alpha_{\rm vir}$ is the virial parameter given by eq. 2, which measures the relative importance of turbulent and thermal pressure relative to gravitational force. The characteristic values of $\epsilon_{\rm ff}\approx 0.01-0.1$ measured in the simulations thus correspond to the characteristic values of $\alpha_{\rm vir}\approx 3-10$, which are set by the dynamics of the ISM gas and stellar feedback that affects the thermal and turbulent pressure of the gas.
Large-scale ISM flows lead to compression and expansion of the gas and fluctuations of $\alpha_{\rm vir}$ (e.g., Semenov et al., 2017). As a local compression decreases $\alpha_{\rm vir}$ locally, $\epsilon_{\mathrm{ff}}$ grows as $\propto\exp(-\sqrt{\alpha_{\rm vir}})$. This growth is accompanied by the growth of star formation rate and associated stellar feedback, which increases local thermal and, importantly, turbulent pressure. This limits the decrease of $\alpha_{\rm vir}$ and leads to self-regulation of its values to the characteristic range of $\alpha_{\rm vir}\approx 3-10$. The characteristic values of $\epsilon_{\rm ff}\approx 0.01-0.1$ in our simulations are thus the result of the star formation prescription and the stellar feedback model adopted in our simulations.
This also allows us to interpret the universality of $\epsilon_{\mathrm{ff}}$ in environments of different metallicity, UV flux, and densities demonstrated above. The pressure and $\alpha_{\rm vir}$ in star-forming regions in these simulations are dominated by turbulent contribution and turbulent pressure. The turbulence is driven by the dynamics of the gas (mainly compression and expansion) and the dissipation of supersonic turbulence, which dominates in star-forming regions, is set by the local turbulent crossing time and does not depend on the metallicity, UV flux, or density. In contrast, the gas cooling and heating, and therefore thermal pressure, do depend on these properties and thus, when $\alpha_{\rm vir}$ is dominated by the thermal pressure, the efficiency per free-fall time varies more strongly with the properties of the environment. This is illustrated explicitly in the Appendix A, where we show $\epsilon_{\mathrm{ff}}$ as a function of metallicity, UV flux, and surface density in simulations, in which subgrid turbulence is not modeled and its contribution to $\alpha_{\rm vir}$ is ignored. The turbulence thus plays an important role in making $\epsilon_{\mathrm{ff}}$ value range universal.
@@
We note, however, that comparing detailed distribution of $\epsilon_{\mathrm{ff}}$ estimated in observations and simulations is not straightforward. In simulations, instantaneous values of $\epsilon_{\mathrm{ff}}$ are considered, while in observations, these values are estimated using instantaneous gas properties and star formation rate indicators that are sensitive to star formation over a certain time scale. The gas density is expected to evolve strongly as the star formation and feedback proceed in a given region (Feldmann & Gnedin, 2011). The large apparent scatter of observational estimates of $\epsilon_{\mathrm{ff}}$ thus largely reflects diverse evolutionary stages of the regions in which these estimates are made (e.g., Feldmann & Gnedin, 2011; Lee et al., 2016; Grisdale et al., 2019). Nevertheless, the simulations shed the light why the characteristic average values of $\epsilon_{\mathrm{ff}}$ do not vary with environment properties.
@@ 5 Summary and conclusions
    We show that the star formation efficiency per free-fall time in $\approx 10$ pc star-forming regions of the simulated disks has characteristic values in the range $\epsilon_{\rm ff}\approx 0.01-0.1$ (Figures 1 and 2). We show that the characteristic values of $\epsilon_{\mathrm{ff}}$ do not vary with metallicity and exhibit only very weak trend with UV flux (Figures 3 and 4).
@@
    Likewise, we show that $\epsilon_{\mathrm{ff}}$ estimated on the $500$ pc scale does not vary with metallicity and shows only weak trends with average UV flux and gas surface density estimated at this scale (Figures 5 and 6).
@@
    We show explicitly that $\epsilon_{\rm ff}$ values are do depend on the metallicity and UV flux in simulations where turbulent pressure contribution to the virial parameter is ignored (Figure 7).
@@ Acknowledgements.
A.K. was supported by the National Science Foundation grant AST-1911111 and NASA ATP grant 80NSSC20K0512. V.S. is grateful for the support provided by Harvard University through the Institute for Theory and Computation Fellowship. The simulations used in this work were carried out on the Midway cluster maintained by the University of Chicago Research Computing Center. Analyses presented in this paper were greatly aided by the following free software packages: NumPy (Harris et al., 2020), SciPy (Jones et al., 2001–), Matplotlib (Hunter, 2007), AstroPy (Astropy Collaboration et al., 2013, 2018, 2022) and yt (Turk et al., 2011). We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.
@@ Appendix A Turbulence-free star formation
All other analysis is done in the same way as in Section 3.1, and we do not change any of the physics associated with star formation or feedback within the simulation. For direct comparison between Figure 5 and 7, we make the same cut of $\alpha_{\mathrm{vir}}\leq 15$, with cells weighted by their star formation rates.

### v2

@@ 1 Introduction
On kiloparsec and larger scales galaxies also form stars inefficiently, which is partly related to the local efficiency per free-fall time (e.g., Krumholz et al., 2012; Semenov et al., 2018). The connection between efficiency in star-forming regions and star formation rate measured on large scales arises because the local inefficiency of star-forming regions and their short lifetimes result in only a small fraction of gas being converted into stars in each star-forming region. Therefore, a given gas parcel has to go through a large number of star formation-dispersal cycles, each taking $\approx 50-100$ Myr as ISM atoms make their way into a star-forming region (Semenov et al., 2017).
@@
[…]
Figure 1: The gas mass-weighted $\sigma_{\mathrm{tot}}-n$ phase space for four of our simulation runs. The color indicates gas temperature (see side color bar). The figure shows gas with the highest star formation rate (roughly $\alpha_{\rm vir}\lesssim 15$) has temperatures (10-30 K) and densities ($10-1000\,\rm cm^{-3}$), which is similar to the typical temperatures and densities of observed star-forming giant molecular clouds when averaged on similar scales. As discussed in Polzin et al. (2024), the low metallicity runs have a lower star formation rate, not because $\epsilon_{\mathrm{ff}}$ is intrinsically lower in low $Z$ gas, but because the amount of cold, dense gas decreases.
@@ 2 Simulations
We use a suite of simulations of an isolated disk galaxy adopting initial conditions similar to the properties of the NGC 300 galaxy (Semenov et al., 2021). These simulations are run using the ART N-body$+$hydrodynamics code (Kravtsov, 1999; Kravtsov et al., 2002; Rudd et al., 2008; Gnedin & Kravtsov, 2011) and include self-consistent modeling of radiative transfer (RT; Gnedin 2014), non-equilibrium chemistry of hydrogen, helium, and molecular hydrogen, as well as realistic prescriptions for star formation and feedback. The latter is sufficiently realistic to replicate detailed properties of star forming regions in NGC 300 (Semenov et al., 2021), and the resolution ($\Delta x=10$  pc in the highest refinement level grid cells, where most star formation is occurring) is sufficient to model feedback-regulated density fluctuations in the simulated interstellar medium.
@@
where the total sub-grid velocity dispersion, $\sigma_{\mathrm{tot}}=\sqrt{\sigma_{\mathrm{t}}^{2}+c_{\mathrm{s}}^{2}}$, includes contributions of both turbulence and the thermal energy of the gas. Note that the virial parameter depends on both the gas temperature through the sound speed and on the turbulent velocity dispersion. This, for example, prevents star formation in the warm gas ($T\gtrsim 10^{3}$ K). Note also that there is no explicit dependence of $\epsilon_{\rm ff}$ on the Mach number because no significant dependence on the Mach number was found in simulations by Padoan et al. (2012) and more recent simulations by Brucy et al. (2024). Figure 1 shows the $\sigma_{\mathrm{tot}}-n$ phase space in our simulations. The dashed line shows the constant value of $\alpha_{\mathrm{vir}}=15$, below which most of the star formation occurs. The figure shows that with the prescription we use in our simulations stars form in cold gas ($T\sim 10-30$ K) with densities of $n\sim 10-1000^{3}\rm\,cm^{-3}$ similar to the densities of observed star-forming molecular clouds averaged on similar scales (tens of parsecs). It also shows that the star-forming gas is different than the gas selected above a constant density and or below a given temperature threshold.
The local value of $\epsilon_{\rm ff}$ in simulation cells evolves with the evolution of the interstellar medium, based on its density and the level of turbulence on small scales. The latter is modeled using an explicit subgrid turbulence model (Semenov et al., 2016).
@@ 3.1 Star formation efficiency as a function of metallicity on $\sim 10$ pc scale
We note these characteristic efficiency values are not a result of tuning but arise naturally in simulations (see Section 4 for discussion). Furthermore, the top panels in Figures 4 and Figure 5 show that the characteristic $\epsilon_{\mathrm{ff}}$ range does not depend on gas metallicity and depends only very weakly on free-space [^1] Free-space UV flux is the flux at a given location not attenuated by local extinction. Krumholz et al. (2009), for example, use free-space flux to mean the flux incident on molecular clouds. In our simulations, the radiative transfer calculations do not include absorption by H2 lines and thus do not model the radiation field self-consistently inside molecular-rich regions and have to rely on the subgrid model. In this case, the free-space flux is the flux returned by the radiative transfer solver and has the physical meaning of the incident field on the molecular gas. UV flux, $U_{\mathrm{MW}}$, normalized by the Milky Way value at 1000 Å (Draine, 1978; Mathis et al., 1983).
Remarkably, while the amount of star-forming gas decreases with decreasing metallicity due to increasingly inefficient cooling (see Polzin et al., 2024, for a detailed discussion), the range of $\epsilon_{\rm ff}$ remains the same for all simulated metallicities. In what follows we show that $\epsilon_{\rm ff}$ also has no, or very weak, dependence on the averaging scale and other properties of the local galactic environment.
@@
We note that the width of the star-formation weighted rate distribution shown in Figures 4 and Figure 5 is comparable to the $\approx 0.3$ dex scatter of $\epsilon_{\rm ff}$ estimated by Hu et al. (2022) for observed star-forming clouds. However, previous studies indicated that the evolution of star-forming cloud properties can contribute significantly to the apparent scatter of $\epsilon_{\rm ff}$ estimated in observations. For example, Grisdale et al. (2019) show that dramatic changes in the properties of star-forming clouds during their evolution contribute most of the scatter in $\epsilon_{\rm ff}$ values estimated for star-forming molecular clouds the way it is done in observations. Notably, they show that the distribution of $\epsilon_{\rm ff}$ estimated in observations by Lee et al. (2016) can be matched in a simulation in which intrinsic $\epsilon_{\rm ff}$ in star-forming molecular gas is fixed to a constant value.
@@ 3.2 Star formation efficiency as a function of local environment properties on 500 pc scale
We examine the star formation efficiency per free-fall time ($\epsilon_{\mathrm{ff}}$), the free-fall time ($t_{\mathrm{ff}}$), and the depletion time ($t_{\mathrm{ff}}/\epsilon_{\mathrm{ff}}$), averaged in patches of 500 pc size, as a function of metallicity, UV flux, and baryon (stars and gas) surface density. Namely, we compute the weighted medians [^2] We define weighted quantiles by the value corresponding to the percentile of the cumulative distribution of weights ordered by the product of the values and the weights. of these quantities using the normalized local star formation rate in each cell as a weight. We also only use the cells with $\alpha_{\mathrm{vir}}\leq 15$, which contribute the bulk of the total star formation in each simulated galaxy. Note that we actually compute the SFR-weighted averages of $1/t_{\mathrm{ff}}$ and $\epsilon_{\mathrm{ff}}/t_{\mathrm{ff}}$ because the star formation rate is proportional to these quantities. However, we plot their inverses $t_{\mathrm{ff}}$ and $t_{\mathrm{ff}}/\epsilon_{\mathrm{ff}}$ since the free-fall time and the depletion time values are more easily interpretable.
@@
Also, over the entire range, the median values of the efficiency are in the range $\approx 0.005-0.05$. Note that this range of values is a bit different than the range of $\epsilon_{\mathrm{ff}}$ in the $\epsilon_{\rm ff}-Z$ panel quoted above because averaging in bins of $U_{\rm MW}$ is over somewhat different gas elements than averaging in bins of $Z$.
@@ 4 Discussion
As noted above, the characteristic values of $\epsilon_{\mathrm{ff}}$in our simulations arise naturally from the specific star-formation prescription based on the simulations of star formation in turbulent molecular clouds (see eq. 1 above; Padoan et al., 2012). The efficiency in this simulation-calibrated prescription is an exponential function of $\sqrt{\alpha_{\rm vir}}$, where $\alpha_{\rm vir}$ is the virial parameter given by eq. 2, which measures the relative importance of turbulent and thermal pressure relative to gravitational force.
The characteristic values of $\epsilon_{\rm ff}\approx 0.01-0.1$ measured in the simulations thus correspond to the characteristic values of $\alpha_{\rm vir}\approx 3-10$, which are set by the dynamics of the ISM gas and stellar feedback that affects the thermal and turbulent pressure of the gas.
Large-scale ISM flows lead to compression and expansion of the gas and fluctuations of $\alpha_{\rm vir}$ (e.g., Semenov et al., 2017). As a local compression decreases $\alpha_{\rm vir}$ locally, $\epsilon_{\mathrm{ff}}$ grows as $\propto\exp(-\sqrt{\alpha_{\rm vir}})$. This growth is accompanied by the growth of star formation rate and associated stellar feedback, which increases local thermal and, importantly, turbulent pressure. This limits the decrease of $\alpha_{\rm vir}$ and leads to self-regulation of its values to the characteristic range of $\alpha_{\rm vir}\approx 3-10$. Part of the reason behind the narrow range of $\alpha_{\rm vir}$ is the fact that during compression and expansion due to large-scale flows and feedback turbulence often evolves adiabatically, $\sigma\propto n^{1/3}$, and the virial parameter changes only mildly during such evolution, $\alpha_{\rm vir}\propto\sigma^{2}/n\propto n^{-1/3}$. The characteristic values of $\epsilon_{\rm ff}\approx 0.01-0.1$ in our simulations are thus the result of the star formation prescription and the stellar feedback model adopted in our simulations.
This also allows us to interpret the universality of $\epsilon_{\mathrm{ff}}$ in environments of different metallicity, UV flux, and densities demonstrated above. The pressure and $\alpha_{\rm vir}$ in star-forming regions in these simulations are dominated by turbulence. The turbulent energy is driven by gas compression by large-scale flows and the dissipation on the turbulent crossing time scale. Both processes do not depend on the metallicity, UV flux, or density. In contrast, the gas cooling and heating, and therefore thermal pressure, do depend on these properties and thus, when $\alpha_{\rm vir}$ is dominated by the thermal pressure, the efficiency per free-fall time varies more strongly with the properties of the environment. This is illustrated explicitly in the Appendix A, where we show $\epsilon_{\mathrm{ff}}$ as a function of metallicity, UV flux, and surface density in simulations, in which subgrid turbulence is not modeled and its contribution to $\alpha_{\rm vir}$ is ignored. The turbulence thus plays an important role in making $\epsilon_{\mathrm{ff}}$ value range universal.
@@
We note, however, that comparing detailed distribution of $\epsilon_{\mathrm{ff}}$ estimated in observations and simulations is not straightforward. In simulations, instantaneous values of $\epsilon_{\mathrm{ff}}$ are considered, while in observations, these values are estimated using instantaneous gas properties and star formation rate indicators that are sensitive to star formation over a certain time scale. The gas density is expected to evolve strongly as the star formation and feedback proceed in a given region (Feldmann & Gnedin, 2011). The large apparent scatter of observational estimates of $\epsilon_{\mathrm{ff}}$ thus largely reflects diverse evolutionary stages of the regions in which these estimates are made (e.g., Feldmann & Gnedin, 2011; Lee et al., 2016; Grisdale et al., 2019). Nevertheless, the simulations shed light on why the characteristic average values of $\epsilon_{\mathrm{ff}}$ do not vary with environment properties.
Finally, we note that recently Semenov et al. (2024) and Segovia Otero et al. (2024) presented results of cosmological galaxy formation simulations of Milky Way-sized galaxies with the same prescription for $\epsilon_{\mathrm{ff}}$ given by equations 1 and 2 above. In both studies, the authors found that the average $\epsilon_{\mathrm{ff}}$ of the simulated galaxies was $\sim 1\%$, although rapidly changing conditions during early stages of evolution result in significant efficiency fluctuations around this characteristic value. In particular, Semenov et al. (2024) showed that fluctuations of $\epsilon_{\mathrm{ff}}$ are the dominant source of burstiness of star formation during the early stages of galaxy evolution.
@@ 5 Summary and conclusions
    We show that the star formation efficiency per free-fall time in $\approx 10$ pc star-forming regions of the simulated disks has characteristic values in the range $\epsilon_{\rm ff}\approx 0.01-0.1$ (Figures 2 and 3). We show that the characteristic values of $\epsilon_{\mathrm{ff}}$ do not vary with metallicity and exhibit only a very weak trend with UV flux (Figures 4 and 5).
@@
    Likewise, we show that $\epsilon_{\mathrm{ff}}$ estimated on a $500$ pc scale does not vary with metallicity and shows only weak trends with average UV flux and gas surface density estimated at this scale (Figures 6 and 7).
@@
    We show explicitly that $\epsilon_{\rm ff}$ values do depend on the metallicity and UV flux in simulations where turbulent pressure contribution to the virial parameter is ignored (Figure 8).
@@ Acknowledgements.
The authors thank the referees for comments that improved the manuscript. A.K. was supported by the National Science Foundation grant AST-1911111 and NASA ATP grant 80NSSC20K0512. V.S. is grateful for the support provided by Harvard University through the Institute for Theory and Computation Fellowship. The simulations used in this work were carried out on the Midway cluster maintained by the University of Chicago Research Computing Center. Analyses presented in this paper were greatly aided by the following free software packages: NumPy (Harris et al., 2020), SciPy (Jones et al., 2001), Matplotlib (Hunter, 2007), AstroPy (Astropy Collaboration et al., 2013, 2018, 2022) and yt (Turk et al., 2011). We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.
@@ Appendix A Turbulence-free star formation
All other analysis is done in the same way as in Section 3.1, and we do not change any of the physics associated with star formation or feedback within the simulation. For direct comparison between Figure 6 and 8, we make the same cut of $\alpha_{\mathrm{vir}}\leq 15$, with cells weighted by their star formation rates.
@@
Note that stellar feedback still operates in the runs without turbulence. Thus, as gas cools and reaches higher densities star formation rate increases and so does associated stellar feedback. The feedback prevents runaway collapse and disperses star-forming regions.